The reaction for producing glucose in plants, called photosynthesis, is

6 CO2 + 6 H2O → (light) C6H12O6 + 6 O2

If a plant produces 2.51 mol C6H12O6, how many moles of H2O are needed?
1
Expert's answer
2019-10-08T07:40:31-0400

6 CO2 + 6 H2O → (light) C6H12O6 + 6 O2

n(C6H12O6)=2.51 mol


n(H2O)=6*n(C6H12O6)=6*2.51 mol=15.06 mol